# On visualisation of completions of free groups

### v1

@@ 1. Introduction
1.1. Free Lie groups and pronilpotent completions of free groups. Denote by $F_{n}$ the free group with $n$ generators. Let $F_{n}^{(k)}$ be its lower central series. According a general construction of Malcev [10] (see, also its exposition in [6], Sect. 17), any finitely generated discrete nilpotent group without torsion canonically embeds as a cocompact lattice to a nilpotent Lie group. In particular, we can apply this construction to the group $F_{n}(k):=F_{n}/F_{n}^{(k)}$ and get a nilpotent Lie group, say $\mathrm{Fr}_{n}(k)$. This Lie group corresponds to the real Lie algebra $\mathfrak{fr}_{n}{(k)}$ with generators $x_{1}$, …, $x_{n}$ and relations $[x_{i_{1}},[x_{i_{2}},\dots,[x_{i_{k}},x_{i_{k+1}}]\dots]]=0$ for all tuples $x_{i_{1}}$, …, $x_{i_{k+1}}$.
@@
so $\alpha$ ranges in the set $\mathcal{A}$ of ordered collections $(\alpha_{1},\alpha_{2},\dots,\alpha_{k})$, where $k=0$, $1$, $2$, $\dots$ and $1\leqslant\alpha_{j}\leqslant n$. We write elements of $\mathrm{Afr}_{n}(\mathcal{E})$ as formal series
@@
2.3. The free Lie algebra and free Lie group. Consider the free Lie algebra $\mathfrak{fr}_{n}(\mathcal{E})$ generated by $\omega_{1}$, …, $\omega_{n}$. For any multiple commutator of elements $\omega_{j}$ we assign a polynomial expression in $\omega_{1}$, …, $\omega_{n}$, assigning to any subcommutator $[x,y]$ the associative expression $xy-yx$. For instance,
$[\omega_{1},\omega_{2}]\mapsto\omega_{1}\omega_{2}-\omega_{2}\omega_{1},\quad[[\omega_{1},\omega_{2}],\omega_{5}]\mapsto(\omega_{1}\omega_{2}-\omega_{2}\omega_{1})\omega_{5}-\omega_{5}(\omega_{1}\omega_{2}-\omega_{2}\omega_{1}),$
@@
[…]
@@ 3. The closure of the free group
Consider a Lyndon word $u$ of length $>1$ and its longest suffix $w$ that is a Lyndon word [^4] One letter is a Lyndon word, so the set of Lyndon suffixes is nonempty.. Then the corresponding prefix $v$ is also a Lyndon word (see [8], Proposition 5.1.3). We write $u=(v\circ w)$ with brackets ()($\quad$)( ), after this we repeat the same reasoning for Lyndon words $v$, $w$, etc. So, we get a parenthesization of the word $u$ (see, e.g., [8]), something like
@@
b) Let $p$ be a prime number. Let $\mathcal{E}$ be the $p$-adic field ${\mathbb{Q}}_{p}$, so ¯$\overline{\mathbb{Z}}={\mathbb{Z}}_{p}$ is the ring of $p$-adic integers.
@@
c) Consider the adelic sum $\mathbb{A}$ of all $p$-adic fields ${\mathbb{Q}}_{p}$, i.e., we consider the product ${\mathbb{Q}}_{2}\times{\mathbb{Q}}_{3}\times{\mathbb{Q}}_{5}\times\dots$ and the subring $\mathbb{A}$ in this product consisting of sequences $(z_{2},z_{3},z_{5},z_{7},\dots)$ such that $|z_{p}|\leqslant 1$ for all but a finite number of primes. We consider the diagonal embedding ${\mathbb{Z}}\to\mathbb{A}$, i.e., m($m\mapsto(m,m,m,\dots)$. Then ¯$\overline{\mathbb{Z}}=\prod_{p}{\mathbb{Z}}_{p}$.
@@ 4. Additional details
We must explain a sense of such product. To evaluate (4.1), it is sufficient to evaluate the product modulo $\mathrm{Afr}_{n}^{[>m]}$ for each $m$. But for all Lyndon words $L$ of length $>m$, we have

### v2

@@ 1. Introduction
1.1. Free Lie groups and pronilpotent completions of free groups. Denote by $F_{n}$ the free group with $n$ generators. Let $F_{n}^{(k)}$ be its lower central series. According a general construction of Malcev [11] (see, also its exposition in [7], Sect. 17), any finitely generated discrete nilpotent group without torsion canonically embeds as a cocompact lattice to a nilpotent Lie group. In particular, we can apply this construction to the group $F_{n}(k):=F_{n}/F_{n}^{(k)}$ and get a nilpotent Lie group, say $\mathrm{Fr}_{n}(k)$. This Lie group corresponds to the real Lie algebra $\mathfrak{fr}_{n}{(k)}$ with generators $x_{1}$, …, $x_{n}$ and relations $x_{i_{1}},[x_{i_{2}},\dots,[x_{i_{k}},x_{i_{k+1}}]\dots]]=0$ for all tuples $x_{i_{1}}$, …, $x_{i_{k+1}}$.
@@
so $\alpha$ ranges in the set $\mathcal{A}$ of ordered collections $(\alpha_{1},\alpha_{2},\dots,\alpha_{k})$, where $k=0$, 11, 22, $\dots$ and $1\leqslant\alpha_{j}\leqslant n$. We write elements of $\mathrm{Afr}_{n}(\mathcal{E})$ as formal series
@@
2.3. The free Lie algebra and free Lie group. Consider the free Lie algebra $\mathfrak{fr}_{n}(\mathcal{E})$ generated by $\omega_{1}$, …, $\omega_{n}$. For any multiple commutator of elements $\omega_{j}$ we assign a polynomial expression in $\omega_{1}$, …, $\omega_{n}$, assigning to any subcommutator $[x,y]$ the associative expression xyyxxy-yx. For instance,
$\omega_{1},\omega_{2}]\mapsto\omega_{1}\omega_{2}-\omega_{2}\omega_{1},\quad[[\omega_{1},\omega_{2}],\omega_{5}]\mapsto(\omega_{1}\omega_{2}-\omega_{2}\omega_{1})\omega_{5}-\omega_{5}(\omega_{1}\omega_{2}-\omega_{2}\omega_{1}),$
@@
$\delta(\omega_{j}^{m})=\sum_{\alpha,\beta:\,0\leqslant\alpha\leqslant m,\,0\leqslant\beta\leqslant m,\,\alpha+\beta\geqslant m}\frac{m!}{(m-\alpha)!\,(m-\beta)!\,(\alpha+\beta-m)!}\,\omega_{j}^{\alpha}\otimes\omega_{j}^{\beta}.$
@@ 3. The closure of the free group
Consider a Lyndon word $u$ of length >1>1 and its longest suffix $w$ that is a Lyndon word [^4] One letter is a Lyndon word, so the set of Lyndon suffixes is nonempty.. Then the corresponding prefix $v$ is also a Lyndon word (see [9], Proposition 5.1.3). We write $u=(v\circ w)$ with brackets $(\quad)$, after this we repeat the same reasoning for Lyndon words $v$, $w$, etc. So, we get a parenthesization of the word $u$ (see, e.g., [9]), something like
@@
b) Let $p$ be a prime number. Let $\mathcal{E}$ be the $p$-adic field ${\mathbb{Q}}_{p}$, so $\overline{{\mathbb{Z}}}={\mathbb{Z}}_{p}$ is the ring of $p$-adic integers. As a corollary, we get the following statements (it is contained in Furusho, Enriquez [3], Lemma 6.3).
@@
c) Consider the adelic sum $\mathbb{A}$ of all $p$-adic fields ${\mathbb{Q}}_{p}$, i.e., we consider the product ${\mathbb{Q}}_{2}\times{\mathbb{Q}}_{3}\times{\mathbb{Q}}_{5}\times\dots$ and the subring $\mathbb{A}$ in this product consisting of sequences $(z_{2},z_{3},z_{5},z_{7},\dots)$ such that $|z_{p}|\leqslant 1$ for all but a finite number of primes. We consider the diagonal embedding ${\mathbb{Z}}\to\mathbb{A}$, i.e., $m\mapsto(m,m,m,\dots)$. Then $\overline{{\mathbb{Z}}}=\prod_{p}{\mathbb{Z}}_{p}$.
@@ 4. Additional details
We must explain a sense of such product. To evaluate (4.1), it is sufficient to evaluate the product modulo $\mathrm{Afr}_{n}^{[>m]}$ for each $m$. But for all Lyndon words $L$ of length >m>m, we have